===== Areas below… =====
* Annex work
* Digital reader
* Wiki work
* Blog work
* Data Farm
* Millennium automation
* Linux install
===== Annex work… =====
* Manually deleted old page-slips. (11/14)
* Emailed Ann answering question (she’s on vacation this week). (11/16)
* Emailed Ann Bonnie’s question, whether there is a limitation on field lengths. (11/16)
* Emailed Bonnie problem with count issue. (11/16)
* Discovered problem with Annex software not processing my auto-generated files; spent *lots* of time investigating. (11/17)
* Emailed Paul auto-parsed data for manual Annex processing. (11/17 through 11/20)
* Emailed Chris & Ann detailed description of problem and investigation. (11/17)
* Ann stopped and restarted the Annex batch process and emailed me that test-files were then processed fine. (11/20)
* Ann emailed me back about field-length question. (11/20)
* Followup with Ann on adding Josiah item # to Annex Packing List.
* Followup with Ann re email notification script.
* Tweak telnet code; improve error-messages.
* Clean up Josiah-to-Annex telnet code.
* Consider auto-deletion of printed page-slips.
===== Digital reader… =====
* No work on this this week. (Picard weblog work relevant to Patrick.)
* Add to configuration page…
* Mets URL root
* JP2 imageserver URL
* three image-size references (done last week; will remove)
* Follow up with Tomcat upgrade when dl-server is upgraded.
* Add thumbnail display capability.
===== Wiki work… =====
* Worked with Patrick on wiki-syntax for displaying images in a table. (11/15)
* Got Tatyana her requested password. (11/16)
* Continue wiki gardening.
* Look into LDAP authentication
===== Blogs/RSS for Steve E. Thompson group… =====
* Emailed Kathryn Rotondo Steve’s requested changes. (11/16)
* Spoke with Kathryn re whether it’s possible to know what users are authorized to use a blog from the admin interface. She didn’t think so. (11/16)
* Emailed Kathryn re problem with one user’s username/password combo. (11/17)
* Emailed Steve that blog is ready, suggested possible next steps. (11/17)
* See if interface allows me to know what users are authorized to use the blog. (done; will remove)
* Figure out how the name of the blog can be changed if necessary.
* Ask folk if they want the email ‘notifications’ feature.
* Install Opera for Steve as a possible RSS reader for notifications.
===== Data Farm / Web-log analysis… =====
* Emailed Bob Morse answering some questions. (11/14)
* Emailed Nancy Magers that her picard log-rotation is working. (11/14)
* Bob Morse emailed noting he had set up a cron job to display AWStats of our pages. (11/14)
* Showed Jean of demo of picard logs viewed through AWStats. (11/14)
* Showed Dan demo of picard logs viewed through AWStats. (11/14)
* Thanked Bob for his AWStats work and suggested possibilities for raw-log transfer. (11/16)
* Began work on weblog-summary page. (11/18)
* Transfer plan from birkinWiki page to wsWiki/weblog-analysis page. That plan has detailed steps. Next few are below.
* Followup with Bob Morse re webpub logs after Thanksgiving.
* Ensure picard logs are working. (done; will remove)
* Perform manual temporary captures while working on automated captures. (no longer necessary; will remove)
* Implement auto captures, merges, and presentation. (in progress)
* Review other Universities’ sites.
* Determine rough schedule with Bonnie to visit/analyze individuals’ workflows.
===== Millennium automation =====
* Bonnied noted possible macro-need from Sam. (11/16)
* Implement the more complex macro.
* Update wiki.
* Let noblenet.org autoit-clearinghouse folk and possibly iii userlist know of work.
* Consider WS Hecker presentation on tool.
===== Linux install =====
* More printing experimentation. (11/16 & 17)
* Configure printing to HP 1200. (in progress)
* Look into Yvonne’s Arc software.
* If IP changes again, consider program emailing IP changes.
===== Other… =====
* Showed AJAX ethereal sample at staff meeting. (11/14)
* Attended NEASIS&T Web conference. (11/15)
* Pursuing getting voicemail. (11/16 & 17)
* Made GUI-search wiki page public and emailed Bonnie and Jean. (11/17)